OpenStudy (anonymous):

1/y-4 + 1/6 =-4/3y-12

OpenStudy (anonymous):

this is fractions

hero (hero):

Yes, I know

OpenStudy (anonymous):

can you help me you think?

hero (hero):

Yes, but you're going to have to learn how to post fractions if you want proper help

OpenStudy (anonymous):

how do i post fractions?

hero (hero):

Well, if you're going to post the linear form, you have to use parentheses to separate the numerator from the denominator

OpenStudy (anonymous):

use my equation as a example so i know?

hero (hero):

For example: (1)/(y-4) + (1)/(6) = (-4)/(3y-12)

hero (hero):

I know it may seem like a lot, but it helps avoid confusion later

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

\[\frac{1}{y-4} + \frac{1}{6} = \frac{-4}{3y-12}\] This is done by typing "\frac{1}{y-4} + \frac{1}{6} = \frac{-4}{3y-12}" without the quotes in the box that pops up when you click "Equation"

hero (hero):

sometimes people just post them like this: 1/(y - 4) + 1/6 = -4/(3y-12) and sometimes that's good enough

hero (hero):

Or you can do what alexray just demonstrated to be even clearer

hero (hero):

Actually, if you use the equation editor you can simply type frac{1}{y - 4} + frac{1}{6} = frac{-4}{3y-12} and avoid all of the extra detail